OpenStudy (anonymous):

If x^2-63x-64=0 and p and n are integers such that p^n=x which of the following CANNOT be a value of p ? a) -8 b) -4 c) -1 d) 4 e) 64

OpenStudy (anonymous):

Use the general solution to the quadratic to find the 2-member solution set. Those will be your values for x (do one at a time for the following): Plug in a) through e) one at a time for p and see if you can find an integer n so that p^n equals first the first value of x and then try the second. You will quickly start eliminating choices.

OpenStudy (anonymous):

Start by solving the quadratic equation for the 2 values of x.

OpenStudy (anonymous):

How did you get that

OpenStudy (phi):

@leah Did you factor the quadratic? what did you get?

OpenStudy (anonymous):

Thats the part im stuck on

OpenStudy (phi):

to factor, 1st look at the -64. the minus means the roots are different signs list the factors of 64 1,64 2,32 4,16 8,8 which pair, when added with different signs, gives -63 (from -63x term)

OpenStudy (anonymous):

1 and 64

OpenStudy (anonymous):

No, not 1 and 64, but -1 and 64.

OpenStudy (phi):

specifically +1 and -64. this means the factors are (x+1)(x-64)= 0 (use FOIL to check this matches x^2 - 63x -64 ) we now know either x+1 is zero or x-64 is zero. x+1=0 or x= -1 and x-64= 0 or x= +64

OpenStudy (phi):

but to answer the question, x is either -1 or +64 let's try the values for p a) p= -8 can you find an integer n so that (-8)^2 = 64 (or -1)

OpenStudy (phi):

-1 and 64 added together is +63. you want -63. see post above

OpenStudy (phi):

a) p= -8 can you find an integer n so that (-8)^2 = 64

OpenStudy (phi):

n=2 make -8^n = -*^2= 64 p= -4: -4^3= -4*-4*-4= -64 this looks bad p= -1: -1^1 = -1 matches -1 can you finish the last two?

OpenStudy (anonymous):

Where did you get n=2

OpenStudy (anonymous):

Never mind i got it

OpenStudy (phi):

the question asks p and n are integers such that p^n=x with p= -8 (choice (a)) I know from the multiplication table that -8*-8= 64 so n=2 (square -8)

OpenStudy (anonymous):

Yes i get it thank you